OpenStudy (anonymous):

Help please! Carmen can retile the kitchen floor in 8 h, but if Joe helps her, the job will take only 3.5 h. How long would it take Joe to do the job alone?

hartnn (hartnn):

if A can do the work in 'x' hrs, and B can do the work in 'y' hrs, then together they will do the work in xy/(x+y) hrs. here xy/(x+y) = 3.5 x=8 find y

OpenStudy (anonymous):

I got 6/2/9 or 6.22222.... Is that right?

hartnn (hartnn):

lets check 8y = 3.5 (8+y) 8y = 14 + 3.5 y 4.5 y = 14 y= 14/4.5 = 28 /9 = ... hrs

OpenStudy (anonymous):

3.5(8) = 28

hartnn (hartnn):

sorry 8y = 3.5 (8+y) 8y = 28 + 3.5 y 4.5 y =28 y= 28/4.5 = 56/9 = ... hrs

OpenStudy (anonymous):

Oh thanks :D i got the answer right

hartnn (hartnn):

good! welcome ^_^